the work function of iron metal is 4.50 eV. what is the maximum speed of the photoelectrons ejected from iron, when light with a wavelength of 142 nm is shined on its surface?
Answer #1

E = hc -34 8 6162720x3y10 & -18 1.393 lo 142x10-9 J je 3 ggx60-16 - ey J. GX 18-19 8.7 hey Krb- E-work function K.E- 8.74er-uanera haux.6x10 - 12/31 11 / 2 (5.) X10-31). v² = 6.24 x26x10-19 v=1:22 x 10 m/sec Drywer